F G QUESTION 3 Dagbreek CC produces three products: AB and C. The following table gives detail of the cost per unit for each of the three products Direct Material: Direct Labour: Variable Manufacturing Overheads3 Total Contribution per Unit Selling Price per Unit 27.00 12.00 3.00 42.00 18.00 60.00 14.00 40.00 32.00 16.00 8.00 4.00 54.00 60.00 36.00 20.00 90.00 80.00 10 12 13 The direct labour rate is R 8 per hour and only 3,000 labour hours per week are available. The marketing division feels that the maximum sales in units per week are 1,000 units 2,000 units 400 units 15 16 17 18 19 20 21 22 a Management wants to know which product should be produced to ensure maximum profit 23 b Management also wants to know the amount of such profits 24 C You have a meeting with the CEO of Dagbreek. Explain to him the difference between absorption costing & variable costing Fixed manufacturing overheads amount to R 20,000 per week Required 25
Answer #1

a. 1000 units of Product A and 111units Of Product B and 400 units of Product C should be produced to ensure maximum profit

Particulars

Product A

Product B

Product C

Selling Price per unit ( X)

60

90

80

Direct Material

27

14

40

Direct Labour

12

32

16

Variable Manufacturing Overheads

3

8

4

Total Variable Cost per unit( Y)

42

54

60

Total Contribution per unit

18

36

20

Labour Hours Required per unit

1.5 hours

4.5 hours

2.5 hours

Contribution per labour hour

12/hour

8/hour

8/hour

Ranking on basis of labour hours

1st

2nd

2nd

b. Maximum Profit =R 18000+11996=29996-20000=R 9996 i.e Approximately R 10,000

In Dagbreek CC, the limiting factor that determines which products to be produced is the Labour Hours. Since contribution per labour hour is maximum in case of Product A , we should produce maximum of product A that is thew maximum permissible units which is=1000 units(Contribution from Product A is 1000 units*18/unit= R 18000.

Therefore for 1000 units of A we use 1500 hours(1000 units*1.5 hours). The remaining 1500 hours can be used to produce only Product B or only Product C or a combination of Product B and Product C as the contribution per labour hour is the same for both products B and C.

Therefore the three alternatives are evaluated in the table given below:

Particulars

only Product B

Maximum of Product C

and balance of Product B

Number of units to be produced

1500 hours/4.5

=333.33=333 units

1500hours/2.5

=600 units but maximum is only 400 units so balance hours to

Be used to produce Product B(500/4.5)

=111 UNITS

400 units of Product C & 111units

Of Product B

Contribution

333 units*36/unit

=R 11988

400 units*20/unit+111 units*36/unit

=R 11996

The difference between the alternatives is due to the rounding of for units to be produced

Maximum Profit =R 18000+11996=29996-20000=R 9996 i.e Approximately R 10,000

c. The difference between absorption costing and variable costing is as follows:

1. In absorption costing fixed manufacturing overheads or expenses are treated as product costs and in the case of variable costing the fixed manufacturing overheads or expenses are treated as period costs.

2. . In absorption costing the stock is valued at the total production costs i.e both the fixed and the variable cost are taken into account thereby increasing g the value of closing stock. In the case of variable costing only variable costs are considered for the stock valuation and therefore that leads to a lower value of the closing stock.

3. In absorption costing we determine the results by calculating the net profits whereas in variable costing we calculate the contribution.

4. The absorption costing is used primarily for external reporting purposes and the variable costing is used for internal reporting purposes.